OpenStudy (anonymous):

Ron paid $799.99 per laptop for three laptops and a flat rate fee for antivirus installation for a total bill of $2,529.57 at Gigahertz Computers. Jerry paid $777.16 per laptop for three laptops plus a per laptop antivirus installation fee for a total bill of $2,481.42 at TechZone Computers. What is the difference in the amounts they paid for the antivirus installation charges? $21.34 $20.34 $23.44 $20.24

OpenStudy (anonymous):

@micahm @musicmouse @TheSmartOne

OpenStudy (anonymous):

@StudyGurl14 @insa @thomaster

OpenStudy (studygurl14):

Ron: 799.99(3) + x =2,529.57 Jerry: 777.16(3) + 3y = 2481.42 Solve for x and y in those equations above^ Then find: |x - y| = ?

OpenStudy (studygurl14):

I multiplied the 799.99 and 777.16 by 3 because that is the cost of each laptop, and three were bought. I multiplied y by 3, because it said there was a "per laptop antivirus installation fee".

OpenStudy (studygurl14):

Do you understand? @MelodyMusic

OpenStudy (anonymous):

hmmm. I think i do :) @StudyGurl14 . can you help with more please?

OpenStudy (studygurl14):

Sure @MelodyMusic . Please close this question and open a new one. Tag me to it.

OpenStudy (anonymous):

Okay :)